The sum of three consecutive terms of an A.P. is 18 and their product is 162.Find the terms.
Answer by mananth(16946)

The sum of three consecutive terms of an A.P. is 18 and their product is 162.Find the terms.
let the three terms be
a-d, a , a+d
a-d+a+a+d =18
3a=18
a=6
(a-d)(a+d)(a) = 162
plug a =6
(6-d)(6+d)(6)=162


(36-d^2)*6 =162
36-d^2= 162/6
36-d^2= 27
d^2=9
d= +/-3
when d=3
3,6,9
when d=-3
9,6,3